OpenStudy (anonymous):

The following function defines a recursive sequence: f(0) = -4 f(1) = 12 f(n) = -3•f(n -1) - 2•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? -4, 12, -28, 60, … -4, -12, -28, -60, … -4, 12, -18, 54, … -4, 12, -18, -54, …

jimthompson5910 (jim_thompson5910):

are you able to find f(2) ?

OpenStudy (anonymous):

i dont understand how to figure this question out at all i need alot of help

jimthompson5910 (jim_thompson5910):

You start with f(n) = -3•f(n -1) - 2•f(n - 2)

jimthompson5910 (jim_thompson5910):

and you replace every n with 2 to get f(2) = -3*f(2-1) - 2*f(2-2)

jimthompson5910 (jim_thompson5910):

f(2) = -3*f(2-1) - 2*f(2-2) turns into f(2) = -3*f(1) - 2*f(0) do you see what to do from here?

OpenStudy (anonymous):

is it f(2)= 44!!!! :)

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@jim_thompson5910 ?

jimthompson5910 (jim_thompson5910):

incorrect

jimthompson5910 (jim_thompson5910):

here's what you should get for f(2) f(n) = -3*f(n-1) - 2*f(n-2) f(2) = -3*f(2-1) - 2*f(2-2) f(2) = -3*f(1) - 2*f(0) f(2) = -3*12 - 2*(-4) f(2) = -36 + 8 f(2) = -28

jimthompson5910 (jim_thompson5910):

my guess is that you somehow got 36+8 = 44 but keep in mind that 36 is negative

OpenStudy (anonymous):

okay so now i know its -28 but there are two choices that have it so how can i figure out which of the two it is

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

f(0) = -4 that's the first term

jimthompson5910 (jim_thompson5910):

basically -4 is the first term

jimthompson5910 (jim_thompson5910):

f(1) = 12 so 12 is the second term

jimthompson5910 (jim_thompson5910):

we just found that f(2) = -28 so the third term is -28

jimthompson5910 (jim_thompson5910):

let's find f(3) f(n) = -3*f(n-1) - 2*f(n-2) f(3) = -3*f(3-1) - 2*f(3-2) f(3) = -3*f(2) - 2*f(1) f(3) = -3*(-28) - 2*12 f(3) = 84 - 24 f(3) = 60

jimthompson5910 (jim_thompson5910):

therefore, the fourth term is 60

OpenStudy (anonymous):

oh my gosh thank you so much for the help i have some more questions like this and if you can help ill open up a new one @jim_thompson5910

jimthompson5910 (jim_thompson5910):

sure I can help with a few more

OpenStudy (anonymous):

thanks!!